(b) 10 points Find all complex numbers z satisfying 28 – 324 – 4 = 0....please with good write hand
Answer #1

solution in the Now given equation is 78 - 374 -4 =0 28 - 3 2 - 4 = 28 - (4-1) 74_4 = 78 - 4*4 +7+-4 = 24 (24-4) +1 (24-41 =